OpenStudy (austinl):

Find an equation of the line segment joining the first point to the second point. (0,0,0) and (1,2,3) Any thoughts guys.

OpenStudy (austinl):

The answer is evidently \(r(t)=t \langle1,2,3\rangle\), where \(0\le t\le 1\)

OpenStudy (austinl):

I want to know how to get there. @myininaya Could you lend a hand?

OpenStudy (science0229):

First, you have to find the direction vector of the equation. Do you know how?

OpenStudy (austinl):

Not a clue.

OpenStudy (science0229):

Let's think of the points, (0,0,0), and (1,2,3) as the components of 2 vectors.

OpenStudy (science0229):

Let's say that the vector (0,0,0) is OA, and (1,2,3) is OB. (O=origin)

OpenStudy (science0229):

To find the direction vector, AB, we have to do OB-OA, which is (1,2,3)-(0,0,0), or (1,2,3).

OpenStudy (austinl):

Well that makes more sense... thanks!

OpenStudy (science0229):

Welcome!

OpenStudy (austinl):

What is that interval about though?

OpenStudy (science0229):

Ignore my last reply. By restricting t from 0 to 1, it defines the line from P to Q.

OpenStudy (loser66):

you have a formula to find it out. It says \(r(t) = (1-t) r_0+t(r_1)\) \(0\leq t\leq 1\) your \(r_0 =(0,0,0) ~~and~~r_1=(1,2,3)\) therefore r(t) = t (1,2,3)